What is the slope-intercept form of y-4=-6/x(x-2)?
Answer by TimothyLamb(4379)

[ What is the slope-intercept form of y-4=-6/x(x-2)? ]
---
The slope-intercept form of a linear equation is: y = mx + b
---
However, your equation: y-4=-6/x(x-2)
is not a linear equation, and so it cannot be put into slope-intercept form.
Linear equations cannot contain variables to powers higher than 1.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php